OpenStudy (anonymous):

The volume of a rectangular prism is 183 in³. What is the volume of a pyramid which has the same height and a congruent base? The volume of a rectangular prism is 183 in³. What is the volume of a pyramid which has the same height and a congruent base? @Mathematics

OpenStudy (anonymous):

please show me the steps cuz im lost

OpenStudy (anonymous):

vol of pyramid = 1/3 of vol of rec prism vol of pyramid = 1/3 * 183 = 61